Find two numbers whose sum is 56 and such that one is 3 times as large as the other.
Answer by HyperBrain(694)

Let x=unknown integer
3x=larger integer
x+3x=56
4x=56
x=56/4=14
3x=3*14=42
Therefore, the two numbers are 14 and 42.
Power up,
HyperBrain!
Stop this mess! Click here